OpenStudy (anonymous):

How do i solve these trig identities? (1 + cosx)/sinx = cot(x/2) cos^6x + sin^6x = cos2x(1-(1/4)sin^2 2x) also, how can I work with Eg. Tan4x, sin2^2x etc?

OpenStudy (mertsj):

also, how can I work with Eg. Tan4x, sin2^2x etc? Use double angle formulas and sin(2x) = 2sin(x)cos(x)

OpenStudy (mertsj):

\[\cot (x/2) = 1/\tan (x/2)=1+\cos(x)/\sin(x)\]

OpenStudy (mertsj):

Is the right side of that second identity \[\cos(2x)(1-\frac14\sin ^{2}(2x))\] ?

OpenStudy (anonymous):

thank you

OpenStudy (mertsj):

Could you please answer my question?

OpenStudy (anonymous):

Yes it is

OpenStudy (mertsj):

Are you sure it is cos^6x + sin^6x? Because I get cos^6x - sin^6x

OpenStudy (anonymous):

Oops my bad, it's cos6x - sin^6x

OpenStudy (anonymous):

How do you deal with the 1/4sin^2x?

OpenStudy (mertsj):

I will show you.

OpenStudy (mertsj):

\[\cos (2x)[1-\frac14\sin ^{2}(2x)]\]

OpenStudy (mertsj):

\[(\cos ^{2}x - \sin ^{2}x)[1/\frac14(2\sin x \cos x)^{2}\]

OpenStudy (mertsj):

That should be 1-1/4 of course

OpenStudy (mertsj):

\[(\cos ^{2}x - \sin ^{2}x)[1- \frac14(4\sin ^{2}x \cos ^{2}x)]\]

OpenStudy (mertsj):

\[(\cos ^{2}x - \sin ^{2}x)[1-\sin ^{2}x \cos ^{2}x]\]

OpenStudy (mertsj):

\[(\cos ^{2}x - \sin ^{2}x)[1-\sin ^{2}x(1-\sin ^{2}x)]\]

OpenStudy (mertsj):

\[(\cos ^{2}x - \sin ^{2}x)[1-\sin ^{2}x +\sin ^{4}x]\]

OpenStudy (mertsj):

Now replace 1 with sin^2x + cos^2x

OpenStudy (mertsj):

\[(\cos ^{2}x - \sin ^{2}x)[\sin ^{2}x + \cos ^{2}x - \sin ^{2}x + \sin ^{4}x]\]

OpenStudy (mertsj):

\[(\cos ^{2}x - \sin ^{2}x)[\cos ^{2}x + \sin ^{4}x\]

OpenStudy (mertsj):

now use FOIL and multiply the two binomials

OpenStudy (mertsj):

\[\cos ^{4}x + \cos ^{2}x \sin ^{4}x - \sin ^{2}x \cos ^{2}x - \sin ^{6}x\]

OpenStudy (mertsj):

Group the middle two terms and factor out cos^2xsin^2x

OpenStudy (mertsj):

\[\cos ^{4}x + \cos ^{2}x \sin ^{2}x(\sin ^{2}x - 1)\]

OpenStudy (mertsj):

Now replace the sin^2x inside the parentheses with 1 - cos^2x

OpenStudy (mertsj):

\[\cos ^{4}x+\cos ^{2}x \sin ^{2}x(1-\cos ^{2}x -1)\]

OpenStudy (mertsj):

\[\cos ^{4}x +\cos ^{2}x \sin ^{2}x(-\cos ^{2}x)\]

OpenStudy (mertsj):

\[\cos ^{4}x-\cos ^{4}x \sin ^{2}x - \sin ^{6}\]x

OpenStudy (mertsj):

\[\cos ^{4}x(1-\sin ^{2}x)- \sin ^{6}\]

OpenStudy (mertsj):

\[\cos ^{4}x(\cos ^{2}x)- \sin ^{6}x\]

OpenStudy (mertsj):

You probably figured out that I dropped the sin^6 for a few steps. You should drag it along.

OpenStudy (anonymous):

Thank you very much!

OpenStudy (mertsj):

Do you understand?

OpenStudy (anonymous):

Yes

OpenStudy (mertsj):

That was a hard one!!!